OpenStudy (anonymous):

Can someone check my SAT questions answers : Point slope question and slope intercept. Define a line that passes through (-1,4) and (-3,2) slope 1. y-4=1(x--1) y-4=x-1 y=x+3 Define a line that passes through (3,2) and has a slope of 3 y-2=3(x-3) y-2=3x-9 y=3x+7 Define a line that has a slope of -1 and y intercept of 4 y=mx+b y=-1x+4 What is the slope of a line defined by y=-12x +89 y=mx+b -12

OpenStudy (nincompoop):

there are several methods you can employ, but traditionally we first calculate for the slope then use the point-slope form to solve for the equation

OpenStudy (nincompoop):

\[m = \frac{y_2-y_1}{x_2-x_1}\] \[y-y_1=m(x-x_1)\] solve for m then use the second equation to solve for y

OpenStudy (nincompoop):

show me your solution, so I know where you need help with

OpenStudy (anonymous):

i posted in original quesiton now. thanks for your help, can you help me check

OpenStudy (anonymous):

are they all right?

OpenStudy (nincompoop):

y-4=1(x--1) do that part again

OpenStudy (anonymous):

Define a line that passes through (-1,4) and (-3,2) (2-4)/(-3--1) Slope = 1 y-y1 = m(x-x1) y-4 = 1(x--1) y-4 = x-1 y = x+3

OpenStudy (nincompoop):

y-4=1(x--1) y-4 = x+1

OpenStudy (nincompoop):

negative and negative makes positive

OpenStudy (anonymous):

Oh my gosh! Thanks for your help haha. Silly mistakes.

OpenStudy (nincompoop):

the rest is the same concept

OpenStudy (anonymous):

so the rest are correct?

OpenStudy (nincompoop):

I don't know, that is something you will need to figure out by double checking

OpenStudy (nincompoop):

y = mx + b b is the y-intercept right?

OpenStudy (anonymous):

yes.

OpenStudy (nincompoop):

then you did well

OpenStudy (nincompoop):

next time, you can also just put all in the same equation \[y-y_1 =\frac{y_2-y_1}{x_2-x_1}(x-x_1)\]

OpenStudy (nincompoop):

this kind of technique is a good practice for physics

OpenStudy (anonymous):

alright thanks ill memorize that! haha but would it work for a question like line c passess through (3,6) and the origin and line d is perpendicular to line c. which of the following could be the equation of line d? ah im working it out atm. but how do u know which coordinate is x1 or/y1 if its like that word problem above (0,0), (3,2) or (3,2) (0,0)